OpenStudy (princess123):

Jack is building a rectangular dog pen that he wishes to enclose. The width of the pen is 2 yards less than the length. If the area of the dog pen is 15 square yards, Jack needs how many yards of fencing to completely enclose the pen?

OpenStudy (anonymous):

let l for the length, W for the width A= L*w 15=L*(L-2) 15=L^2-2L 0 = L^2 -2L -15

OpenStudy (anonymous):

use , quadratic equation form solve